# Beyond Pixels: Exploring DOM Downsampling for LLM-Based Web Agents

### v1

@@ 3 DOM Downsampling
Rooted in signal processing, downsampling defines a technique for reducing data that scales out of time or space constraints. Chunks of data are thereby locally consolidated, while assuming relevant information is retained to a high degree O’Shea and Nash (2015). Broadly speaking, a JPEG image stores only an average colour for patches of pixels Wallace (1992). Effects of such downsampling visually increase with patch size, whereas the depicted object keeps being recognisable up to a large patch size333 The cover page conceptualises downsampling through an example image, and analogously for HTML (serialised DOM). . Related concepts are, in fact, implemented behind LLM vision APIs to subsidise image input OpenAI (2025c).
@@ 3.1 Downsampled DOM Snapshots
The DOM structurally resembles a tree. Each node represents a semantic entity. Both node syntax, and individual content may imply a UI feature – a button element with text “Submit” is likely clickable Haine (2007). There are three serialisable types of redundant DOM nodes: element, text, and attribute444 For consistency, our algorithmic design treats attributes as element node child nodes, and furthermore DOM leaf nodes.  MDN Contributors (l). We tie downsampling, i.e., consolidation, to node syntax. The degree to which a node represents a UI feature is not exclusively an objective matter. To reconnect with reported HTML interpretation strengths of LLMs, we draw ground truth about HTML (DOM) semantics from the latest GPT-4o (gpt-4o-2024-11-20OpenAI (c). The ground truth ultimately constitutes a UI feature degree rating per node type. Element nodes are, moreover, sub-classified by high-level purpose. While ratings could be perceived as algorithmic input, we herein imply those as constants. Attachment A lists the complete ground truth, including seminal model prompts.
@@ 3.2 Element Downsampling
if Depth(ELEMENT) % k == 0:
@@ E $\leftarrow$ { ELEMENT, Parent(ELEMENT) }
elements dictate a UI’s text formatting – not to be confused with individual text. Formatting induces nuanced semantics. Consider bold text, where stroke weight emphasises a word. Or a table, which spatially relates words among each other. D2Snap translates content elements to Markdown555 Markdown is a semantic equivalent to a the content element subset of HTML. We respect an extended Markdown flavour that includes, i.a., tables GitHub (2019).  Gruber and Swartz (2004). At that, we utilise its idea of being significantly more comprehensive than HTML. This works, as mixed code interpretation abilities of LLMs were previously supported Pian et al. (2023). Structurally, content translation corresponds to replacing the element by a single text node that contains its Markdown equivalent.
@@ 4.1 Dataset
We ask a third individual to solve each task under supervision. For each notably different UI, we serialise the respective web application state – both GUI, and DOM. We also serialise a grounded GUI snapshot close to how it is done by Browser Use666 We capture grounded GUI snapshots as follows: Visible interactive elements are enhanced with a coloured bounding box and a numerical identifier. The identifier is listed as text, supplemented with element tag name and contained text. We source our script from the Browser Use repository Browser Use (b) (https://github.com/surfly/D2Snap/blob/3788eb5d6f7d056d4a1f22cd100f0eea79d7fc27/snapshots/_highlight.js)  Müller and Žunič (2024). As an example, Attachment D displays a (grounded) GUI snapshot from the evaluation dataset. The seed dataset expands to a total of 52 records. A record, in other words, represents a partial solution space regarding a comprehensive web task.
@@ 4.2 Procedure
To control the effects of different model system prompts, we create a short, generic template OpenAI (d). The system prompt template requests all elements that are required to interact with in order to solve an incidental web task. It is substituted with regard to snapshot class specifics: elements are supposed to be targeted in a format that aligns with the snapshot, such as CSS selectors. Attachment E shows the system prompt template, as well as concrete substitutes per snapshot class. Our evaluation framework iterates through the dataset, and prompts the latest GPT-4o (gpt-4o-2024-11-20OpenAI (b, c) with different subject snapshots, and the respective system prompt. If the set of suggested elements corresponds to a superset of any set in the record- and snapshot-respective reference, we count it a success777 During evaluation, elements are compared with slight tolerance: For GUI snapshots, we inflate the referenced bounding boxes by 10 pixels. For DOM-based snapshots, we as well accept a referenced parent, child or sibling node. .
@@ 5.2 Accessible Representations
Reader views and accessibility trees have been around to enhance human readability of web application UIs. Such representations are retrieved directly from a DOM, but no longer resemble HTML, which disposes of inherent UI interpretation capabilities of LLMs.
@@ 6.1 Contributions
The repository related to this paper is available at https://github.com/surfly/D2Snap [^8] The paper repository contains a D2Snap implementation, the evaluation dataset, and the evaluation framework (including a system prompt template)., [^9] The paper repository is redundantly archived at https://github.com/t-ski/D2Snap..

### v2

@@ 3 DOM Downsampling
Rooted in signal processing, downsampling defines a technique for reducing data that scales out of time or space constraints. Chunks of data are thereby locally consolidated, while assuming relevant information is retained to a high degree O’Shea and Nash (2015). Broadly speaking, a JPEG image stores only an average colour for patches of pixels Wallace (1992). Effects of such downsampling visually increase with patch size, whereas the depicted object keeps being recognisable up to a large patch size [^3] The cover page conceptualises downsampling through an example image, and analogously for HTML (serialised DOM). . Related concepts are, in fact, implemented behind LLM vision APIs to subsidise image input OpenAI (2025c).
@@ 3.1 Downsampled DOM Snapshots
The DOM structurally resembles a tree. Each node represents a semantic entity. Both node syntax, and individual content may imply a UI feature – a button element with text “Submit” is likely clickable Haine (2007). There are three serialisable types of redundant DOM nodes: element, text, and attribute [^4] For consistency, our algorithmic design treats attributes as element node child nodes, and furthermore DOM leaf nodes.  MDN Contributors (l). We tie downsampling, i.e., consolidation, to node syntax. The degree to which a node represents a UI feature is not exclusively an objective matter. To reconnect with reported HTML interpretation strengths of LLMs, we draw ground truth about HTML (DOM) semantics from the latest GPT-4o (gpt-4o-2024-11-20OpenAI (c). The ground truth ultimately constitutes a UI feature degree rating per node type. Element nodes are, moreover, sub-classified by high-level purpose. While ratings could be perceived as algorithmic input, we herein imply those as constants. Attachment A lists the complete ground truth, including seminal model prompts.
@@ 3.2 Element Downsampling
if Depth(ELEMENT) % k == 0
or Class(Parent(ELEMENT)) not container’:
@@ E $\leftarrow$ { ELEMENT, Parent(ELEMENT) }
elements dictate a UI’s text formatting – not to be confused with individual text. Formatting induces nuanced semantics. Consider bold text, where stroke weight emphasises a word. Or a table, which spatially relates words among each other. D2Snap translates content elements to Markdown [^5] Markdown is a semantic equivalent to a the content element subset of HTML. We respect an extended Markdown flavour that includes, i.a., tables GitHub (2019).  Gruber and Swartz (2004). At that, we utilise its idea of being significantly more comprehensive than HTML. This works, as mixed code interpretation abilities of LLMs were previously supported Pian et al. (2023). Structurally, content translation corresponds to replacing the element by a single text node that contains its Markdown equivalent.
@@ 4.1 Dataset
We ask a third individual to solve each task under supervision. For each notably different UI, we serialise the respective web application state – both GUI, and DOM. We also serialise a grounded GUI snapshot close to how it is done by Browser Use [^6] We capture grounded GUI snapshots as follows: Visible interactive elements are enhanced with a coloured bounding box and a numerical identifier. The identifier is listed as text, supplemented with element tag name and contained text. We source our script from the Browser Use repository Browser Use (b) (https://github.com/surfly/D2Snap/blob/3788eb5d6f7d056d4a1f22cd100f0eea79d7fc27/snapshots/_highlight.js)  Müller and Žunič (2024). As an example, Attachment D displays a (grounded) GUI snapshot from the evaluation dataset. The seed dataset expands to a total of 52 records. A record, in other words, represents a partial solution space regarding a comprehensive web task.
@@ 4.2 Procedure
To control the effects of different model system prompts, we create a short, generic template OpenAI (d). The system prompt template requests all elements that are required to interact with in order to solve an incidental web task. It is substituted with regard to snapshot class specifics: elements are supposed to be targeted in a format that aligns with the snapshot, such as CSS selectors. Attachment E shows the system prompt template, as well as concrete substitutes per snapshot class. Our evaluation framework iterates through the dataset, and prompts the latest GPT-4o (gpt-4o-2024-11-20OpenAI (b, c) with different subject snapshots, and the respective system prompt. If the set of suggested elements corresponds to a superset of any set in the record- and snapshot-respective reference, we count it a success [^7] During evaluation, elements are compared with slight tolerance: For GUI snapshots, we inflate the referenced bounding boxes by 10 pixels. For DOM-based snapshots, we as well accept a referenced parent, child or sibling node. .
@@ 5.2 Accessible Representations
Reader views and accessibility trees have existed to enhance readability of web application UIs for humans. Such representations are retrieved directly from a DOM, but no longer resemble HTML, which disposes of inherent LLM UI interpretation capabilities.
@@ 6.1 Contributions
The repository related to this paper is available at https://github.com/webfuse-com/D2Snap [^8] The paper repository contains a D2Snap implementation, the evaluation dataset, and the evaluation framework (including a system prompt template)., [^9] The paper repository is redundantly archived at https://github.com/t-ski/D2Snap..